OpenStudy (lovelyanna):

Given the linear equation x + y = 5, find the slope of its graph. A. –5 B. 5 C. –1 D. 1 Is it C?

OpenStudy (zpupster):

y=mx+b subtract x both sides y=-x+5 what is your m??

OpenStudy (math&ing001):

Write it as y=-x+5 General form y=m*x+b m is the slope

OpenStudy (zpupster):

C yes. good job!

OpenStudy (lovelyanna):

@math&ing001 so C or correct, or not?

OpenStudy (math&ing001):

Yes :)

OpenStudy (lovelyanna):

Ok, thanks :)

OpenStudy (math&ing001):

Anytime